OpenStudy (anonymous):

Prove the lim n->inf of k^2 * (-1)^k

OpenStudy (zarkon):

there is no n in your expression...so limit would just be k^2 * (-1)^k

OpenStudy (anonymous):

my bad. yes as k goes to infinity

OpenStudy (zarkon):

the limit does not exist

OpenStudy (anonymous):

i know it diverges

OpenStudy (anonymous):

cause its going to oscillate to - and + infinity right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

but do you know how to show/prove that?

OpenStudy (zarkon):

can you prove that \[\lim_{k\to\infty}k^2=\infty\]

OpenStudy (anonymous):

yea through the ratio test

OpenStudy (anonymous):

and the limit test.

OpenStudy (anonymous):

but you cant prove the other one through the ratio test because its not a positive series.

OpenStudy (zarkon):

ratio test is for series